OpenStudy (anonymous):

Help Please!!! The constant force F of magnitude 3lb is applied to the block. the force has the same direction as the vector u=3i+4j. Find the work done in the direction of motion if the block moves from P1 (3,1) to P2 (9,3). Distance measured in feet.

OpenStudy (anonymous):

work = force * distance force = 3 lb distance equals the sqrt( (change in x)^2 + (change in y)^2 )

OpenStudy (anonymous):

But i thought work=force * distance * cos of theta

OpenStudy (anonymous):

that is true too. there are just different ways to solve the same thing so the key in this thing is to find the distance between the two points, can you do it?

OpenStudy (anonymous):

yes sqrt (40) right? \[\sqrt{9-3^{2}+3-1^{2}}\]

OpenStudy (anonymous):

the answer in the back says 78/5 ft/lb

OpenStudy (anonymous):

(9-3)^2 right? not 9-3^2?

OpenStudy (anonymous):

yes thats what i meant

OpenStudy (anonymous):

excellent so distance = sqrt(40) force = 3lb and work = force*distance

OpenStudy (anonymous):

\[\sqrt{40} \times 3 =18.9736\]

OpenStudy (anonymous):

but 78/5 is =15.6